OpenStudy (playgames1122):

What is the cost of a $1,200 washing machine after a 1/5 discount of the original price? Explain.

OpenStudy (anonymous):

here's what i'd do in this case 1/5=2/10 2/10=20% 20%=0.20 multiply 1200 by 0.20

OpenStudy (anonymous):

that's what i'd do

OpenStudy (danjs):

If you get a 1/5 discount, you are paying (1 - 1/5) = 4/5 the cost, 4/5 of 1200